how many grams of Na2SO4 would be formed if 0.75 moles of NaOH reacted


1
Expert's answer
2021-02-25T03:44:46-0500

2NaOH + H2SO4 = Na2SO4 + 2H2O

n(NaOH):n(Na2SO4)=2:1

n(Na2SO4)=0.75/2=0.375mol


m(Na2SO4)=0.375*142=53.25g